10.Parental Nutrition Support. A patient is receiving 1500 milliliters of 50% dextrose and 1500 milliliters of 7% amino acid solution; 500 mL of a 20% lipid emulsion. The total kcalories provided should equal. Show Work:

11. A TPN solution provides 750 mL of a 5% amino acid solution. The nurse calculates that the patient is receiving ___________ grams of protein. Show Work:

  1. A liter of a TPN solution contains 500 milliliters of 50% dextrose solution and 500 milliliters of 5% amino acid solution. Determine the daily and energy need of a person who receives 2-liters per day of such a solution. Calculate the average energy intake if the person also receives 500 milliliters of a 20% fat emulsion three times a week. Show Work:
Answer #1

1.dextrose (1500*50/100ml)* 3.4 =2550kcal

amino acid (1500*7/100)*4 =420 kcal

lipid (500*20 =5000kcal

total k cal = 7970 kcal

2.(750*5/100 )*6.25 =23.4375